### v1

# Mills’ constant is irrational

Abstract.
Let $\lfloor x\rfloor$ denote the integer part of $x$. In 1947, Mills constructed a real number $\xi$ greater than $1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$. We define Mills’ constant as the smallest real number $\xi$ satisfying this property. In this article, we determine that Mills’ constant is irrational. Moreover, we also obtain partial results on the transcendency of the constant.
Key words and phrases:
prime-representing constants, Mills’ constant, transcendental number theory
2020 Mathematics Subject Classification:
11J72, 11J81
1. Introduction
Let $\mathbb{N}$ denote the set of all positive integers. Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$. A complex number $\alpha$ is transcendental if $P(\alpha)\neq 0$ for all polynomials $P\in\mathbb{Q}[x]\setminus\{0\}$.
Mills [12] constructed a constant $\xi>1$ such that
[…]
Such $\xi$ is not uniquely determined. In this article, we define Mills’ constant [^1] There are two types of definitions of Mills’ constant. Finch [7] implicitly introduced Mills’ constant as a real number $\xi$ satisfying (1.1). Our definition can be seen in [3]. as the smallest $\xi>1$ satisfying (1.1). It is well-defined from [13, Remark 4.2], which means that Mills’ constant exists. Assuming the Riemann hypothesis, Caldwell and Cheng [3] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately $1.3063778838\cdots$. It is an unsolved problem to determine whether the constant is rational or irrational. The goal of this article is to solve it.
Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi$ greater than $1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$. By [11, Corollary 4] and [13, Remark 4.2], $\xi_{c}$ exists for every integer $c\geq 2$. A real algebraic integer greater than $1$ is a Pisot number if all of its conjugates over $\mathbb{Q}$ except for itself lie in the open unit disc. We obtain the following results.
Theorem 1.1.
For every integer $c\geq 4$, the constant $\xi_{c}$ is transcendental.
Theorem 1.2.
Either the constant $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
For more generalization, let $(c_{k})_{k\in\mathbb{N}}$ be a real sequence with $c_{k}\geq 1$ for every $k\in\mathbb{N}$. We define $C_{k}=c_{1}\cdots c_{k}$ for every $k\in\mathbb{N}$ and
$\mathcal{W}(c_{k})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
An element $\xi\in\mathcal{W}(c_{k})$ is called a prime-representing constant (for short PRC). Wright [15] firstly investigated the geometric structure of $\mathcal{W}(c_{k})$. Especially, by his result [15, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$. Therefore, there exist infinitely many transcendental numbers in $\mathcal{W}(c_{k})$, but we do not get any arithmetic properties of the smallest element in $\mathcal{W}(c_{k})$ from this result. Alkauskas and Dubickas [1, Theorem 1] firstly gave a result on the transcendency of specific PRCs. They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if a sequence $(c_{k})_{k\in\mathbb{N}}$ of real numbers satisfying
[…]
They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental or not. The author and Takeda [13, Theorem 1.3] showed that the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental if a sequence $(c_{k})_{k\in\mathbb{N}}$ of integers satisfying
[…]
They also obtained a result on the algebraic independency of PRCs in [13]. The condition (iii) is essentially needed in these works, but we accomplish replacing (iii) with the condition $c_{k+1}\geq 4$ for infinitely many positive integers $k$ as follows.
Theorem 1.3.
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying
  1. (1)
    $c_{1}\geq 1$;
  2. (2)
    $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$;
  3. (3)
    $c_{k+1}=b$ for infinitely many positive integers $k$.
Then the smallest real number in $\mathcal{W}(c_{k})$ exists. Let $\xi$ be the number, and the following properties are true:
  • if $b\geq 4$, then $\xi$ is transcendental;
  • if $b=3$, then either $\xi$ is transcendental, or $\xi^{C_{m}}$ is a Pisot number of degree $3$ for some $m\in\mathbb{N}$.
Theorem 1.3 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.3 in Section 4. At the last of this section, let us ask several remaining questions.
Question 1.4.
Is $\xi_{2}$ rational or irrational?
Question 1.5.
Does there exist $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$?
If the answer to Question 1.4 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2. In Section 5, we will give assumptions leading to the transcendency of $\xi_{3}$. We will explain the main difficulties of Questions 1.4 and 1.5 in Remarks 3.6 and 4.4, respectively.
2. Strategy of proofs
In this section, we demonstrate a strategy of proofs. We first recall Mills’ construction, that is, we now construct $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
Let $\mathcal{P}$ be the set of all prime numbers, and let $\theta=21/40$. We apply the following theorem given by Baker, Harman, and Pintz [2] [^2] Mills applied the result given by Ingham [8] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$. Thus, the construction of this section is modernly modified. .
Theorem 2.1.
There exists a constant $d_{0}>0$ such that
#$\#([x,x+x^{\theta}]\cap\mathcal{P})\geq\frac{d_{0}x^{\theta}}{\log x}$
for sufficiently large $x>0$.
Let $p_{1}$ be a sufficiently large prime number. By Theorem 2.1 with $x=p_{1}^{3}$, we find $p_{2}\in\mathcal{P}$ such that
$p_{1}^{3}\leq p_{2}\leq p_{1}^{3}+p_{1}^{3\theta}.$
Since $23\theta=63/40<$ 2 and $p_{1}$ is sufficiently large, we see that
$p_{1}^{3}+p_{1}^{3\theta}<p_{1}^{3}+3p_{1}^{2}+3p_{1}=(p_{1}+1)^{3}-1.$
Similarly, by Theorem 2.1 with $x=p_{2}^{3}$, we find $p_{3}\in\mathcal{P}$ such that
$p_{2}^{3}\leq p_{3}\leq p_{2}^{3}+p_{2}^{3\theta}<(p_{2}+1)^{3}-1.$
By iterating this argument, we find a sequence $(p_{k})_{k=1}^{\infty}$ of prime numbers such that
(2.1) $p_{k}^{3}\leq p_{k+1}\leq p_{k}^{3}+p_{k}^{3\theta}<(p_{k}+1)^{3}-1$
for every $k\in\mathbb{N}$. This leads to
$p_{1}^{1/3^{1}}\leq p_{2}^{1/3^{2}}\leq p_{3}^{1/3^{3}}\leq\cdots<(p_{3}+1)^{1/3^{3}}<(p_{2}+1)^{1/3^{2}}<(p_{1}+1)^{1/3^{1}},$
and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}\eqqcolon w$ and $\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w^{\prime}$ exist [^3] Actually, $w=w^{\prime}$ is true, but we do not use this formula.. Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}\leq w^{\prime 3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
We will further give a proof of the following result for sharing the main idea.
Theorem 2.2.
The constant $w$ is irrational.
Remark 2.3.
We actually want to discuss the arithmetic properties of $\xi_{3}(=\min\mathcal{W}(3))$. There is a gap between $w$ and $\xi_{3}$. However, this gap is not essential to prove the irrationality (and transcendency). We can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.1 and 3.2).
To prove Theorem 2.2, we estimate the least distance between $w^{3^{k}}$ and prime numbers.
Lemma 2.4.
There exists $\gamma>0$ such that for every sufficiently large $k$
$|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
Proof.
Since $p_{k+1}=\lfloor w^{3^{k+1}}\rfloor$ for every $k\in\mathbb{N}$, (2.1) leads to
$p_{k}^{3}\leq p_{k+1}\leq w^{3^{k+1}}\leq p_{k}^{3}+p_{k}^{3\theta}+1\leq p_{k}^{3}(1+2p_{k}^{3(\theta-1)})$
for every sufficiently large $k\in\mathbb{N}$. Therefore,
$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}.$
By applying the Maclaurin expansion, $(1+x)^{1/3}=1+O(x)$ holds if $|x|\leq 1$. Thus, since $\theta-1=-19/40<0$, for every sufficiently large $k$, we have
$p_{k}\leq w^{3^{k}}\leq p_{k}\left(1+O(p_{k}^{3(\theta-1)})\right)=p_{k}+O(p_{k}^{3(\theta-1)+1}),$
and hence there exists $K_{0}>0$ such that for every sufficiently large $k$
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}.$
Combining $03\theta-2=63/40-2<$ 0 and the most left-hand side of (2.1), we obtain
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}\leq K_{0}p_{k-1}^{3(3\theta-2)}\leq\cdots\leq K_{0}p_{1}^{3^{k-1}(3\theta-2)},$
which completes the proof of Lemma 2.4. ∎
Lemma 2.4 means that $w^{3^{k}}$ is very close to some prime number. On the other hand, the powers of a rational number are at a distance from integers. Indeed, in 1957, Mahler [9] proved the following result. Here, for every $x\in\mathbb{R}$, we define
$\|x\|=\min\{|x-m|\colon m\in\mathbb{Z}\}.$
Theorem 2.5.
Let $r/s$ be a rational number greater than $1$ and which is not an integer. Let $\epsilon$ be a positive real number. Then, there exists an integer $n_{0}=n_{0}(r/s,\epsilon)$ such that
$\|(r/s)^{n}\|>e^{-\epsilon n}$
for every integer $n\geq n_{0}$.
Proof of Theorem 2.2.
We first assume that $w$ is an integer. Then
$p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3},$
which is a contradiction since $p_{2}$ is a prime number. Therefore, $w$ is not an integer. We secondly assume that $w$ is a rational number. Since $w$ is not an integer, there exist co-prime integers $r$ and $s$ with $r>s\geq 2$ such that $w=r/s$. By Lemma 2.4, there exists $\gamma>0$ such that for every sufficiently large $k$
$\|w^{3^{k}}\|\leq|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
On the other hand, by Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$, for every sufficiently large $k$
$\|w^{3^{k}}\|>e^{-3^{k}\gamma}.$
This is a contradiction. Therefore, $w$ is irrational. ∎
Through the discussion above, we can share the main idea of the proof of Theorem 1.3. We will compare the following two phenomena:
  • (i)
    powers of the smallest PRC are very close to some prime number;
  • (ii)
    powers of almost all real algebraic numbers are at a distance from integers.
For (ii), we need a generalization of Theorem 2.5 to real algebraic numbers. Corvaja and Zannier [4] gave this generalization by applying the $p$-adic Schmidt subspace theorem. Dubickas used their result and showed the following [6, Lemma 6] to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2).
Theorem 2.6.
Let $\alpha>1$ be an algebraic number, and $q$ be a positive integer. Suppose that $0<s_{1}<s_{2}<\cdots$ is a sequence of positive integers. Then, either for some $m\geq 0$ the number $\alpha^{s_{m}}$ is a Pisot number or for each $\epsilon>0$ there exists a positive integer $k_{0}=k_{0}(\epsilon)$ such that
$\|q\alpha^{s_{k}}\|>e^{-\epsilon s_{k}}$
for every $k\geq k_{0}$.
Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients satisfying its leading coefficient is positive. Wagner and Ziegler [14] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
(2.2) $x_{n+1}=P(x_{n})\quad\text{for $n=0,1,2\ldots$, \quad and $x_{n}\to\infty$ as $n\to\infty$.}$
They showed that $\lim_{n\to\infty}x_{n}^{1/d^{n}}=\alpha>1$ exists and it is either irrational or an integer. Further, Dubickas [6] improved their result to the transcendency of $\alpha$ by applying Theorem 2.6. Their methods are very useful not only for the recursive equations (2.2) but also for recursive inequalities like (2.1). In addition, Dubickas presented the following lemma [6, Lemma 8].
Lemma 2.7.
Let $\beta$ be a Pisot number of degree $\ell\geq 2$ with conjugates $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ labelled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. Then, there is a constant $\lambda>0$ depending only on $\beta$ such that
$|\beta_{2}^{n}+\cdots+\beta_{\ell}^{n}|\geq|\beta_{2}|^{n}n^{-\lambda}$
for each sufficiently large integer $n$.
We will apply Lemma 2.7 to the cases in which a Pisot number appears.
3. Lemmas
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. By [13, Theorem 1.2], the set $\mathcal{W}(c_{k})$ is non-empty, closed, and lower bounded. Therefore, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists. We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by [13, Theorem 1.3]. Thus setting $B=\sup_{k\in\mathbb{N}}c_{k}$, we have $B<\infty$. Let $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$ which is an infinite set from (3). Let $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for all $k\in\mathbb{N}$. Let $\theta=21/40$. The following Lemmas 3.1 and 3.2 can be seen in [13, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.
Lemma 3.1.
For all $k\in\mathbb{N}$, we have
(3.1) $p_{k}^{c_{k+1}}\leq p_{k+1}<(p_{k}+1)^{c_{k+1}}-1.$
Proof.
Take any $k\in\mathbb{N}$. By the definition of $\xi$, we see that $p_{k}\leq\xi^{C_{k}}<p_{k}+1$, which implies that $p_{k}^{c_{k+1}}\leq\xi^{C_{k+1}}<(p_{k}+1)^{c_{k+1}}$. We recall that $c_{k+1}$ is a positive integer greater than or equal to $2$, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$. It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number. Therefore, we obtain (3.1). ∎
Lemma 3.2.
There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
(3.2) $p_{k}^{c_{k+1}}\leq p_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}.$
To prove Lemma 3.2, we apply the following result given by Matomäki (see [11, Lemma 9] and [10]).
Theorem 3.3.
There exist positive constants $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
#$\#([n,n+n^{\gamma}]\cap\mathcal{P})\leq d_{1}\frac{n^{\gamma}}{\log n}.$
By applying this theorem, Matomäki [11] constructed an element in $\mathcal{W}(2)$. It is truly astonishing because Mills’ construction tells us that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to concrete an element in $\mathcal{W}(2)$. This problem is strongly connected with Legendre’s conjecture which asserts that for every $n$ there exists a prime number between $n^{2}$ and $(n+1)^{2}$. It is unsolved and believed to be extremely difficult. However, Theorem 3.3 with $\gamma=1/2$ states that there exists a prime number $p\in[n,n+n^{1/2}]$ except for very few intervals $[n,n+n^{1/2}]$. Therefore, Matomäki discovered that we can construct a PRC if the exceptional intervals are very few. We apply Theorem 3.3 to prove the following lemma.
Lemma 3.4.
Let $c$ be a real number in the interval $[2,B]$, and $\epsilon$ be a sufficiently small positive real number. Let $X_{0}$ be a sufficiently large positive parameter depending only on $B$ and $\epsilon$. Let $X$ and $\eta$ be real numbers satisfying $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$. Suppose that
(3.3) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X}$
for some constant $d_{2}>0$. Then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
(3.4) #$\#([q^{c},q^{c}+q^{c-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c-1}}{\log q^{c}}.$
Proof.
Let $R=[X,X+X^{\eta}]\cap\mathcal{P}$. Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$. Indeed, we have
$X^{c}$ $\leq q^{c}<q^{c}+q^{c-1}=q^{c}(1+q^{-1})\leq(X+X^{\eta})^{c}(1+X^{-1})$
$\leq X^{c}(1+X^{\eta-1})^{B}(1+X^{-1})\leq 2X^{c}$
since $\eta\leq 1-\epsilon$ and $X$ is sufficiently large. Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$. By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
(3.5) $\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
$\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
since $X\geq X_{0}$ and $X_{0}$ is sufficiently large. Then by combining (3.3), (3.5), and Theorem 3.3 with $x\coloneqq X^{c}$ and $\gamma\coloneqq\gamma$, there exists a prime number $q\in R$ satisfying (3.4). ∎
Proof of Lemma 3.2.
Assume that for every $k_{0}>0$ there exists $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$ such that
(3.6) $p_{k+1}\notin[p_{k}^{c_{k+1}},p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}].$
We will find $w\in\mathcal{W}(c_{k})$ such that $w<\xi$ by Matomäki’s construction, which will imply a contradiction to the minimality of $\xi$.
We now take $k_{0}$ as a sufficiently large positive parameter. By the assumption, we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.6). By Lemma 3.1, it is false that $p_{k}^{c_{k+1}}>p_{k+1}$, and hence we have
(3.7) $p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}.$
Then, we set $q_{k}=p_{k}$. Theorem 2.1 with $x\coloneqq q_{k}^{c_{k+1}}$ implies that
#$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
By Lemma 3.4 with $X\coloneqq q_{k}^{c_{k+1}}$, $\eta\coloneqq\theta$, and $c\coloneqq c_{k+2}$, there exists $q_{k+1}\in\mathcal{P}$ satisfying
(3.8) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1})$
(3.9) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}}.$
Let us construct a sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying the following properties: for every $m\geq k$, we have
(3.10) $q_{m}^{c_{m+1}}\leq q_{m+1}\leq q_{m}^{c_{m+1}}+q_{m}^{c_{m+1}-1},$
(3.11) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{3}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}},$
where $d_{3}$ is a small positive constant. In the case $m=k$, both (3.10) and (3.11) are true by (3.8) and (3.9). We assume that there exists a sequence $(q_{m})_{m=k}^{M}$ of prime numbers satisfying (3.10) and (3.11) for every $m=k,\ldots,M-1$, where $M\geq k+1$. Then by applying Lemma 3.4 with $X\coloneqq q_{M}^{c_{M+1}}$, $\eta\coloneqq(c_{M+1}-1)/c_{M+1}$, and $c\coloneqq c_{M+2}$, there exists a prime number $q_{M+1}$ satisfying (3.10) and (3.11) with $m=M$. By induction, we obtain an infinite sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying (3.10) and (3.11).
Furthermore, we set $q_{m}=p_{m}$ for every $m=1,2,\ldots,k-1$. By Lemma 3.1 and (3.10), we obtain
$q_{m}^{c_{m+1}}\leq q_{m+1}<(q_{m}+1)^{c_{m+1}}-1$
for every $m\in\mathbb{N}$. Therefore, we have
(3.12) $q_{m}^{1/C_{m}}\leq q_{m+1}^{1/C_{m+1}}<(q_{m+1}+1)^{1/C_{m+1}}<(q_{m}+1)^{1/C_{m}}$
for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=w$ and $\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}=w^{\prime}$ exist. In addition, $w\in\mathcal{W}(c_{k})$ holds by (3.12). On the other hand, combining (3.7) and (3.8),
$\lfloor w^{C_{k+1}}\rfloor=q_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}=\lfloor\xi^{C_{k+1}}\rfloor.$
This is a contradiction since $\xi$ is the smallest element in $\mathcal{W}(c_{k})$. ∎
Lemma 3.5.
There exist $k_{1}=k_{1}(\xi)>0$ and $K_{0}>0$ such that for every integer $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$, we have
(3.13) $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}},$
where $\theta_{b}=1-\theta-1/b>0$. Further, there exists $\gamma>0$ such that for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$
(3.14) $|\xi^{C_{k}}-p_{k}|\leq e^{-\gamma C_{k}}.$
Proof.
Let $k_{0}(\xi)$ be as in Lemma 3.2, and let $k_{1}=k_{1}(\xi)>k_{0}(\xi)$ be a sufficiently large parameter. Since $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for every $k\in\mathbb{N}$, by Lemma 3.2, each $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$ satisfies
(3.15) $p_{k}^{c_{k+1}}\leq p_{k+1}\leq\xi^{C_{k+1}}\leq p_{k+1}+1\leq p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}}.$
Therefore, for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$, we have
(3.16) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{n}(1+p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}=p_{k}+O(p_{k}^{c_{k+1}(\theta-1)+1}).$
We note that $c_{k+1}(\theta-1)+1=c_{k+1}(\theta-1+\frac{1}{b})=-c_{k+1}\theta_{b}<0$ since $c_{k+1}=b$ and $b\geq 3$. Therefore, we get (3.13). Furthermore, Lemma 3.1 leads to
(3.17) $p_{k}^{c_{k+1}}\geq p_{k-1}^{c_{k+1}c_{k}}\geq p_{k-2}^{c_{k+1}c_{k}c_{k-1}}\geq\cdots\geq p_{1}^{C_{k+1}/c_{1}}.$
Combining (3.13) and (3.17), we obtain (3.14). ∎
Remark 3.6.
We can observe the main difficulties of Question 1.4 in (3.16). If $c_{1}=c_{2}=\cdots=2$, then roughly speaking, (3.16) should be
$p_{k}\leq\xi^{2^{k}}<p_{k}+p_{k}^{2(1/2-1)+1}=p_{k}+1.$
It does not suffice to obtain (3.14). Further, by Matomäki’s construction, we do not know which specific prime number $q$ satisfies (3.4). This is also one of the main difficulties.
4. Proof of Theorem 1.3
Lemma 4.1.
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. Let $\xi$ be the smallest element in $\mathcal{W}(c_{k})$. Then the following properties are true:
  • if $b\geq 5$, then $\xi$ is transcendental;
  • if $b=4$, then either $\xi$ is transcendental, or there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$;
  • if $b=3$, then either $\xi$ is transcendental, or there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$ or $3$.
Proof.
Let $k_{1}$ be as in Lemma 3.5. Assume that $\xi$ is a real algebraic number. By the definition of $\mathcal{W}(c_{k})$, $\xi>1$ holds. By Lemma 3.5, there exists $\gamma>0$ such that for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$
$|\xi^{C_{k}}-p_{k}|\leq e^{-\gamma C_{k}}.$
Therefore, Theorem 2.6 with $\alpha\coloneqq\xi$, $q\coloneqq 1$, $s_{k}\coloneqq C_{k}$, and $\epsilon\coloneqq\gamma$ implies that $\xi^{C_{m}}$ is a Pisot number for some $m\in\mathbb{N}$. Let $\beta=\xi^{C_{m}}$, and let $\ell$ be the degree of $\beta$. We now show the following claim, where $\theta_{b}$ is defined in Lemma 3.5.
Claim. We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
If $\ell=1$, then $\beta(=\xi^{C_{m}})$ is a positive integer. Therefore,
$p_{m+1}=\lfloor\xi^{C_{m+1}}\rfloor=(\xi^{C_{m}})^{c_{m+1}}=p_{m}^{c_{m+1}},$
which is a contradiction since $p_{m+1}$ is a prime number and $c_{m+1}\geq 2$. Therefore $\ell\geq 2$.
We next assume that $\ell>(b\theta_{b})^{-1}+1$. Let $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$ labeled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. For every $k\geq m$, since $C_{k}/C_{m}$ is a positive integer, we obtain
$\beta_{1}^{C_{k}/C_{m}}+\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\in\mathbb{Z}.$
Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. By Lemma 3.5,
(4.1) $\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\|=\|\beta_{1}^{C_{k}/C_{m}}\|=\|\xi^{C_{k}}\|\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}.$
By (3.15), we recall that $\xi^{C_{k+1}}\leq p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}}$, to obtain
(4.2) $\xi^{C_{k+1}}\leq p_{k}^{c_{k+1}}(1+2p_{k}^{(\theta-1)c_{k+1}})\leq 2p_{k}^{c_{k+1}}.$
Therefore, combining (4.1), (4.2), and $c_{k+1}=b$, we have
$\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\|\leq K_{1}\xi^{-C_{k+1}\theta_{b}}=K_{1}\beta_{1}^{-b\theta_{b}C_{k}/C_{m}}$
for some constant $K_{1}>0$. By Lemma 2.7 with $n\coloneqq C_{k}/C_{m}$, there exists $\lambda=\lambda(\beta)>0$ such that
$\left\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\right\|=\left|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\right|\geq|\beta_{2}|^{C_{k}/C_{m}}(C_{k}/C_{m})^{-\lambda}$
since $k$ is sufficiently large. By taking $k\to\infty$ on $k\in\mathcal{I}_{b}$, we conclude that $|\beta_{2}|\leq\beta_{1}^{-b\theta_{b}}$. Since $|\beta_{2}|\geq|\beta_{3}|\geq\cdots\geq|\beta_{l}|$ and $\ell>(b\theta_{b})^{-1}+1$, we have $|\beta_{2}\cdots\beta_{\ell}|\leq\beta_{1}^{-(\ell-1)b\theta_{b}}<\beta_{1}^{-1}$. This is a contradiction to $|\beta_{1}\cdots\beta_{\ell}|\geq 1$. Thus, we obtain the claim.
Case ($b\geq 5$): It follows that
$b\theta_{b}\geq\frac{19\cdot 5}{40}-1\geq\frac{19}{8}-1=\frac{11}{8}\quad\therefore\quad(b\theta_{b})^{-1}+1\leq\frac{8}{11}+1<2,$
which is a contradiction to the claim. Therefore, $\xi$ is transcendental.
Case ($b=4$): By simple calculation,
$b\theta_{b}=\frac{19\cdot 4}{40}-1=\frac{19}{10}-1=\frac{9}{10}\quad\therefore\quad(b\theta_{b})^{-1}+1=\frac{10}{9}+1<3.$
Case ($b=3$): We also observe that
$b\theta_{b}=\frac{19\cdot 3}{40}-1=\frac{57}{40}-1=\frac{17}{40}\quad\therefore\quad\ (b\theta_{b})^{-1}+1=\frac{40}{17}+1<4.$
This completes the proof. ∎
Let us discuss the case when $\xi^{C_{m}}$ is a Pisot number of degree $2$ for some $m\in\mathbb{N}$.
Lemma 4.2.
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. Suppose that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$. We set
$t_{k}=\beta_{1}^{C_{k}/C_{m}}+\cdots+\beta_{\ell}^{C_{k}/C_{m}}\in\mathbb{Z}$
for every $k\geq m$. Then we have $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$ for every sufficiently large $k\in\mathbb{N}$.
Proof.
Let $\delta_{k}=\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}$ for every $k\geq m$. Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. Since $t_{k}$ and $\beta_{1}^{C_{k}/C_{m}}$ are real numbers, $\delta_{k}$ is also a real number. Further, we recall that $|\beta_{j}|<1$ for every $2\leq j\leq\ell$, which implies that $\delta_{k}\in(-1/2,1/2)$ since $k$ is sufficiently large. If $-1/2<\delta_{k}\leq 0$, then $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$. If $20<\delta_{k}<1/$ 2, then by (3.13), $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}$ holds for some constant $K_{0}>0$. Therefore,
$\delta_{k}\leq t_{k}-p_{k}\leq\delta_{k}+K_{0}p_{k}^{-c_{k+1}\theta_{b}}.$
Since $t_{k}-p_{k}\in\mathbb{Z}$ and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough. ∎
Lemma 4.3.
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
Proof.
Assume that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$. Let $\beta=\xi^{C_{m}}$, and let $\beta_{1}=\beta,\beta_{2}$ be the conjugates of $\beta$. Let $t_{k}$ be as in Lemma 4.2.
Case ($b=3$): For every $k\in\mathcal{I}_{3}$, we set $x_{1}=\beta_{1}^{C_{k}/C_{m}}$ and $x_{2}=\beta_{2}^{C_{k}/C_{m}}$. Then we obtain $c_{k+1}=3$ and
$t_{k}^{3}=(x_{1}+x_{2})^{3}=x_{1}^{3}+3x_{1}^{2}x_{2}+3x_{1}x_{2}^{2}+x_{2}^{3}=t_{k+1}+3x_{1}x_{2}t_{k}.$
By applying Lemma 4.2, for sufficiently large $k\in\mathcal{I}_{3}$, we have $p_{k}^{3}=p_{k+1}+3x_{1}x_{2}p_{k}$. By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$. This is a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
Case ($b=4$): Let $k$ be a sufficiently large integer in $\mathcal{I}_{4}\cap[m,\infty)$. Since $\mathcal{I}_{4}$ is infinite, we may assume that $C_{k}/C_{m}(\eqqcolon d_{k})$ is a positive even integer. Setting $x_{1}=\beta_{1}^{d_{k}}$ and $x_{2}=\beta_{2}^{d_{k}}$, we have
$\beta_{1}^{d_{k}}+\beta_{2}^{d_{k}}=p_{k}=\lfloor\beta_{1}^{d_{k}}\rfloor\leq\beta_{1}^{d_{k}}.$
Then $\beta_{2}^{d_{k}}\leq 0$. This is a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
Combining Lemmas 4.1 and 4.3 completes the proof of Theorem 1.3.
Remark 4.4.
To prove the transcendency of Mills’ constant, it remains to show that there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $3$ when $b=3$. In this remark, we describe the difficulty of this case.
Suppose $b=3$, and we assume that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $3$. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\beta_{3}$ be the conjugates of $\beta$. Similarly with the proof of Lemma 4.3, we obtain
(4.3) $p_{k+1}=p_{k}^{3}-3b_{k}p_{k}+3e_{k}$
for every sufficiently large $k\in\mathcal{I}_{3}$, where
$b_{k}\coloneqq(\beta_{1}\beta_{2})^{C_{k}/C_{m}}+(\beta_{2}\beta_{3})^{C_{k}/C_{m}}+(\beta_{3}\beta_{1})^{C_{k}/C_{m}},\quad e_{k}\coloneqq(\beta_{1}\beta_{2}\beta_{3})^{C_{k}/C_{m}}.$
Therefore, we have to investigate properties of $b_{k}$ and $e_{k}$. The author does not have any good ideas on how to treat $b_{k}$ and $e_{k}$ simultaneously. Indeed, the method of the proof of Lemma 4.3 does not work.
5. Further discussions
We will give several assumptions leading to the transcendency of $\xi_{3}$.
Proposition 5.1.
Assume that $\lfloor\xi_{3}^{3^{k}}\rfloor\not\equiv\lfloor\xi_{3}^{3^{k+1}}\rfloor\mod 3$ for infinitely many positive integers $k$. Then $\xi_{3}$ is transcendental.
Proof.
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. Let $p_{k}=\lfloor\xi_{3}^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$. By Fermat’s little theorem, $p_{k}^{3}\equiv p_{k}\mod 3$. Thus, by Remark 4.4, we have $p_{k+1}\equiv p_{k}\mod 3$ for every sufficiently large $k\geq m$. This is a contradiction. ∎
A Mills prime is a prime number of the form $\lfloor\xi_{3}^{3^{k}}\rfloor$. By Proposition 5.1, if there are infinitely many Mills primes of the forms $13n+$ 1 and $23n+$ 2 respectively, then $\xi_{3}$ is transcendental. However, it seems to be impossible to classify Mills primes modulo $3$.
Proposition 5.2.
Assume that there exist $\theta^{\prime}\in(0,1/2)$ and $x_{0}>0$ such that for every $x\geq x_{0}$ we can find a prime number $p\in[x,x+x^{\theta^{\prime}}]$. Then $\xi_{3}$ is transcendental.
Proof.
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. We set $\theta_{b}^{\prime}=\theta_{3}^{\prime}=1-\theta^{\prime}-1/3$. In a similar manner with Lemma 3.5 and the claim in the proof of Lemma 4.1, we have
$3=\deg(\xi_{3}^{3^{m}})\leq(3\theta_{3}^{\prime})^{-1}+1=\frac{1}{2-3\theta^{\prime}}+1<\frac{1}{2-3/2}+1=3,$
which is a contradiction. ∎
Let $p_{n}$ be the $n$th prime number for every $n\in\mathbb{N}$. Cramér [5] conjectured that $p_{n+1}-p_{n}=O((\log p_{n})^{2})$. If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive constant. Therefore, assuming Cramér’s conjecture, $\xi_{3}$ is transcendental.
Acknowledgement
The author would like to thank Professor Shigeki Akiyama for daily attractive discussions on mathematics. The author is grateful to Professor Hajime Kaneko for giving useful ideas on the cases $b=3$ and $4$. The author is also grateful to Professor Wataru Takeda for finding errors and mistakes. The author was supported by JSPS KAKENHI Grant Numbers JP22J00025 and JP22KJ0375.

### v2

@@
 Let $\lfloor x\rfloor$ denote the integer part of $x$.
-In 1947, Mills constructed a real number $\xi$ greater than $1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$.
+In 1947, Mills constructed a real number $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$.
 We define Mills’ constant as the smallest real number $\xi$ satisfying this property.
-In this article, we determine that Mills’ constant is irrational.
-Moreover, we also obtain partial results on the transcendency of the constant.
+Determining whether this number is irrational has been a long-standing problem.
+In this paper, we show that Mills’ constant is irrational.
+Furthermore, we obtain partial results on the transcendency of this number.
 Key words and phrases:
@@
 11J72, 11J81
+It will appear in Mathematika.
 1. Introduction
-Let $\mathbb{N}$ denote the set of all positive integers.
-Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$.
-A complex number $\alpha$ is transcendental if $P(\alpha)\neq 0$ for all polynomials $P\in\mathbb{Q}[x]\setminus\{0\}$.
-Mills [12] constructed a constant $\xi>1$ such that
+Let $\mathbb{N}$ denote the set of all positive integers. Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$. Mills [16] constructed a real number $\xi>1$ such that
 […]
 Such $\xi$ is not uniquely determined.
-In this article, we define Mills’ constant [^1] There are two types of definitions of Mills’ constant.
-Finch [7] implicitly introduced Mills’ constant as a real number $\xi$ satisfying (1.1).
-Our definition can be seen in [3].
-as the smallest $\xi>1$ satisfying (1.1).
-It is well-defined from [13, Remark 4.2], which means that Mills’ constant exists.
-Assuming the Riemann hypothesis, Caldwell and Cheng [3] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately $1.3063778838\cdots$.
-It is an unsolved problem to determine whether the constant is rational or irrational.
-The goal of this article is to solve it.
-Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi$ greater than $1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
-By [11, Corollary 4] and [13, Remark 4.2], $\xi_{c}$ exists for every integer $c\geq 2$.
+In this paper, we define Mills’ constant [10, p.130] as the smallest $\xi>1$ satisfying (1.1).
+In Lemma 3.1, we will show the existence of the smallest element of such $\xi$’s.
+Assuming the Riemann hypothesis, Caldwell and Cheng [4] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately equal to $1.3063778838\cdots$.
+Proving the irrationality of a specific number satisfying (1.1) has remained an unsolved problem for over 70 years since Mills’ result.
+In this paper, we determine that Mills’ constant is irrational.
+Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi>1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
+The number $\xi_{c}$ exists for every integer $c\geq 2$ by Corollary 3.4 (or [15, Corollary 4] and [18, Remark 4.2]).
 A real algebraic integer greater than $1$ is a Pisot number if all of its conjugates over $\mathbb{Q}$ except for itself lie in the open unit disc.
@@
 Theorem 1.1.
-For every integer $c\geq 4$, the constant $\xi_{c}$ is transcendental.
+For every integer $c\geq 4$, $\xi_{c}$ is transcendental.
 Theorem 1.2.
-Either the constant $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
-For more generalization, let $(c_{k})_{k\in\mathbb{N}}$ be a real sequence with $c_{k}\geq 1$ for every $k\in\mathbb{N}$. We define $C_{k}=c_{1}\cdots c_{k}$ for every $k\in\mathbb{N}$ and
-$\mathcal{W}(c_{k})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
+Either $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
+For more generalization, let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers with $c_{k}\geq 1$ for all $k\in\mathbb{N}$, and we define
+$\mathcal{W}(c_{k})=\mathcal{W}((c_{k})_{k=1}^{\infty})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\},$
+where $C_{k}=c_{1}\cdots c_{k}$ for all $k\in\mathbb{N}$.
+Remark 1.3.
+We always use the notation $\mathcal{W}(c_{k})$ as $\mathcal{W}((c_{k})_{k=1}^{\infty})$ for simplicity, and so we remark that $\mathcal{W}(c_{k})$ is independent of $k$.
+For instance, if $(c_{k})_{k=1}^{\infty}=(3,3,3,\ldots)$, then
+$\mathcal{W}(3)=\{\xi>1\colon\lfloor\xi^{3^{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
 An element $\xi\in\mathcal{W}(c_{k})$ is called a prime-representing constant (for short PRC).
-Wright [15] firstly investigated the geometric structure of $\mathcal{W}(c_{k})$.
-Especially, by his result [15, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$.
+After Mills’ result [16], there are several studies for extending this result.
+For example, in 1950, Kuıpers [12] constructed an element in $\mathcal{W}(c)$ for all integers $c\geq 3$.
+In 1951, Ansari [2] extended the range of $c$ to all real numbers $c>77/29=2.655\cdots$.
+In the same period, Niven [17] independently gave a similar extension, but he extended $c$ to real numbers $c>8/3=2.666\cdots$.
+Furthermore, in 1954, Wright [20] first studied the set of PRCs [^1] In honor of Wright, the author uses the symbol “$\mathcal{W}$” as the set of prime-representing constants.
+Wright actually considered a wider class of representing functions..
+We refer the readers to Dudley’s survey [9] for more details.
+It is a good survey of the early research of Mills-type constants.
+Especially, by Wright’s result [20, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$.
 Therefore, there exist infinitely many transcendental numbers in $\mathcal{W}(c_{k})$, but we do not get any arithmetic properties of the smallest element in $\mathcal{W}(c_{k})$ from this result.
-Alkauskas and Dubickas [1, Theorem 1] firstly gave a result on the transcendency of specific PRCs.
-They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if a sequence $(c_{k})_{k\in\mathbb{N}}$ of real numbers satisfying
+Alkauskas and Dubickas [1, Theorem 1] gave a result on the transcendency of specific PRCs.
+They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if the sequence $(c_{k})_{k=1}^{\infty}$ satisfies
 […]
-They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental or not.
-The author and Takeda [13, Theorem 1.3] showed that the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental if a sequence $(c_{k})_{k\in\mathbb{N}}$ of integers satisfying
+They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental. The author and Takeda [18, Theorem 1.3] showed the following result.
+Theorem 1.4.
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of positive integers satisfying
+  1. (i)
+    $c_{1}\geq 1$;
+  2. (ii)
+    $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$;
+  3. (iii)
 […]
-They also obtained a result on the algebraic independency of PRCs in [13].
+Then, the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental.
+The author and Takeda also obtained a result on the algebraic independency of PRCs in [18].
 The condition (iii) is essentially needed in these works, but we accomplish replacing (iii) with the condition $c_{k+1}\geq 4$ for infinitely many positive integers $k$ as follows.
@@
     $c_{k+1}=b$ for infinitely many positive integers $k$.
-Then the smallest real number in $\mathcal{W}(c_{k})$ exists. Let $\xi$ be the number, and the following properties are true:
+Then the smallest real number in $\mathcal{W}(c_{k})$ exists, say $\xi$. For this $\xi$ the following is true:
   • if $b\geq 4$, then $\xi$ is transcendental;
   • if $b=3$, then either $\xi$ is transcendental, or $\xi^{C_{m}}$ is a Pisot number of degree $3$ for some $m\in\mathbb{N}$.
-Theorem 1.3 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.3 in Section 4. At the last of this section, let us ask several remaining questions.
-Question 1.4.
+Theorem 1.5 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.5 in Section 4. At the end of this section, let us ask several remaining questions.
+Question 1.6.
 Is $\xi_{2}$ rational or irrational?
-Question 1.5.
+Question 1.7.
 Does there exist $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$?
-If the answer to Question 1.4 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2.
+If the answer to Question 1.7 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2.
 In Section 5, we will give assumptions leading to the transcendency of $\xi_{3}$.
-We will explain the main difficulties of Questions 1.4 and 1.5 in Remarks 3.6 and 4.4, respectively.
-2. Strategy of proofs
-In this section, we demonstrate a strategy of proofs.
+We will explain the main difficulty of Question 1.7 in Remark 4.4.
+2. Strategy of our proof
+In this section, we demonstrate a strategy of our proof.
 We first recall Mills’ construction, that is, we now construct $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
 Let $\mathcal{P}$ be the set of all prime numbers, and let $\theta=21/40$.
-We apply the following theorem given by Baker, Harman, and Pintz [2] [^2] Mills applied the result given by Ingham [8] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$.
+We apply the following theorem given by Baker, Harman, and Pintz [3, p.562] [^2] Mills applied the result given by Ingham [11] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$.
 Thus, the construction of this section is modernly modified.
@@
 Theorem 2.1.
-There exists a constant $d_{0}>0$ such that
+There exists a real number $d_{0}>0$ such that
 #$\#([x,x+x^{\theta}]\cap\mathcal{P})\geq\frac{d_{0}x^{\theta}}{\log x}$
@@
 $p_{1}^{3}\leq p_{2}\leq p_{1}^{3}+p_{1}^{3\theta}.$
-Since $23\theta=63/40<$ 2 and $p_{1}$ is sufficiently large, we see that
+By $23\theta=63/40<$ 2, we see that
 $p_{1}^{3}+p_{1}^{3\theta}<p_{1}^{3}+3p_{1}^{2}+3p_{1}=(p_{1}+1)^{3}-1.$
@@
 $p_{1}^{1/3^{1}}\leq p_{2}^{1/3^{2}}\leq p_{3}^{1/3^{3}}\leq\cdots<(p_{3}+1)^{1/3^{3}}<(p_{2}+1)^{1/3^{2}}<(p_{1}+1)^{1/3^{1}},$
-and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}\eqqcolon w$ and $\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w^{\prime}$ exist [^3] Actually, $w=w^{\prime}$ is true, but we do not use this formula..
-Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}\leq w^{\prime 3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
-We will further give a proof of the following result for sharing the main idea.
+and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}=\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w$ exists.
+Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
+We will further show the following result for giving a sketch of the proof.
 Theorem 2.2.
-The constant $w$ is irrational.
+The number $w$ is irrational.
 Remark 2.3.
 We actually want to discuss the arithmetic properties of $\xi_{3}(=\min\mathcal{W}(3))$.
-There is a gap between $w$ and $\xi_{3}$.
-However, this gap is not essential to prove the irrationality (and transcendency).
-We can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.1 and 3.2).
-To prove Theorem 2.2, we estimate the least distance between $w^{3^{k}}$ and prime numbers.
+There is a gap between $w$ and $\xi_{3}$, but we can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.5 and 3.6).
+To prove Theorem 2.2, we estimate the distance between $w^{3^{k}}$ and the nearest integer.
 Lemma 2.4.
@@
 $p_{k}^{3}\leq p_{k+1}\leq w^{3^{k+1}}\leq p_{k}^{3}+p_{k}^{3\theta}+1\leq p_{k}^{3}(1+2p_{k}^{3(\theta-1)})$
-for every sufficiently large $k\in\mathbb{N}$. Therefore,
-$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}.$
-By applying the Maclaurin expansion, $(1+x)^{1/3}=1+O(x)$ holds if $|x|\leq 1$. Thus, since $\theta-1=-19/40<0$, for every sufficiently large $k$, we have
-$p_{k}\leq w^{3^{k}}\leq p_{k}\left(1+O(p_{k}^{3(\theta-1)})\right)=p_{k}+O(p_{k}^{3(\theta-1)+1}),$
-and hence there exists $K_{0}>0$ such that for every sufficiently large $k$
-$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}.$
-Combining $03\theta-2=63/40-2<$ 0 and the most left-hand side of (2.1), we obtain
-$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}\leq K_{0}p_{k-1}^{3(3\theta-2)}\leq\cdots\leq K_{0}p_{1}^{3^{k-1}(3\theta-2)},$
-which completes the proof of Lemma 2.4. ∎
-Lemma 2.4 means that $w^{3^{k}}$ is very close to some prime number.
+for every $k\in\mathbb{N}$. Therefore,
+$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}<p_{k}+2p_{k}^{3\theta-2},$
+where the last inequality follows from $(1+x)^{1/3}<1+x$ for all $x>0$. Thus, for every $k\in\mathbb{N}$, we have
+$|w^{3^{k}}-p_{k}|\leq 2p_{k}^{3\theta-2}.$
+Combining $03\theta-2=63/40-2<$ 0 and the first inequality of (2.1), we obtain
+$|w^{3^{k}}-p_{k}|\leq 2p_{k}^{3\theta-2}\leq 2p_{k-1}^{3(3\theta-2)}\leq\cdots\leq 2p_{1}^{3^{k-1}(3\theta-2)},$
+which completes the proof of Lemma 2.4 by selecting $\gamma=\frac{2-3\theta}{6}\log p_{1}>0$. ∎
+Lemma 2.4 means that $w^{3^{k}}$ is very close to some integer (especially, prime number).
 On the other hand, the powers of a rational number are at a distance from integers.
@@
 Theorem 2.5.
-Let $r/s$ be a rational number greater than $1$ and which is not an integer. Let $\epsilon$ be a positive real number. Then, there exists an integer $n_{0}=n_{0}(r/s,\epsilon)$ such that
-$\|(r/s)^{n}\|>e^{-\epsilon n}$
+Let $\alpha$ be a rational number strictly greater than $1$ and which is not an integer.
+Let $\epsilon$ be a positive real number.
+Then, there exists an integer $n_{0}=n_{0}(\alpha,\epsilon)$ such that
+$\|\alpha^{n}\|>e^{-\epsilon n}$
 for every integer $n\geq n_{0}$.
 Proof of Theorem 2.2.
-We first assume that $w$ is an integer. Then
-$p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3},$
-which is a contradiction since $p_{2}$ is a prime number.
-Therefore, $w$ is not an integer.
-We secondly assume that $w$ is a rational number.
-Since $w$ is not an integer, there exist co-prime integers $r$ and $s$ with $r>s\geq 2$ such that $w=r/s$.
-By Lemma 2.4, there exists $\gamma>0$ such that for every sufficiently large $k$
-$\|w^{3^{k}}\|\leq|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
-On the other hand, by Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$, for every sufficiently large $k$
-$\|w^{3^{k}}\|>e^{-3^{k}\gamma}.$
-This is a contradiction. Therefore, $w$ is irrational. ∎
-Through the discussion above, we can share the main idea of the proof of Theorem 1.3. We will compare the following two phenomena:
-  • (i)
-    powers of the smallest PRC are very close to some prime number;
-  • (ii)
+It is clear that $w\notin\mathbb{Z}$ since if not, then we have $p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3}$, a contradiction.
+Therefore, $w$ is irrational by combining Lemma 2.4 and Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$.
+∎
+Through the discussion above, we can give a sketch of the proof of Theorem 1.5. We will compare the following two phenomena:
+  • powers of the smallest PRC are very close to some integer;
   • powers of almost all real algebraic numbers are at a distance from integers.
-For (ii), we need a generalization of Theorem 2.5 to real algebraic numbers.
+For observing the second phenomenon, we need a generalization of Theorem 2.5 to real algebraic numbers.
 Corvaja and Zannier [6] gave this generalization by applying the $p$-adic Schmidt subspace theorem.
-Dubickas used their result and showed the following [6, Lemma 6] to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2).
+Dubickas used their result and prepared the following theorem to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2) (see [8, Lemma 6]).
 Theorem 2.6.
@@
 for every $k\geq k_{0}$.
-Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients satisfying its leading coefficient is positive. Wagner and Ziegler [14] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
+Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients such that its leading coefficient is positive. Wagner and Ziegler [19] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
 (2.2) $x_{n+1}=P(x_{n})\quad\text{for $n=0,1,2\ldots$, \quad and $x_{n}\to\infty$ as $n\to\infty$.}$
@@
 Further, Dubickas [8] improved their result to the transcendency of $\alpha$ by applying Theorem 2.6.
-Their methods are very useful not only for the recursive equations (2.2) but also for recursive inequalities like (2.1).
-In addition, Dubickas presented the following lemma [6, Lemma 8].
+Their methods are instrumental not only for the recursive equations (2.2) but also for recursive inequalities like (2.1).
+In addition, Dubickas presented the following lemma [8, Lemma 8].
 Lemma 2.7.
 Let $\beta$ be a Pisot number of degree $\ell\geq 2$ with conjugates $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ labelled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$.
-Then, there is a constant $\lambda>0$ depending only on $\beta$ such that
+Then, there is a real number $\lambda>0$ depending only on $\beta$ such that
 $|\beta_{2}^{n}+\cdots+\beta_{\ell}^{n}|\geq|\beta_{2}|^{n}n^{-\lambda}$
@@
 We will apply Lemma 2.7 to the cases in which a Pisot number appears.
-3. Lemmas
+3. Lemmas and auxiliary results
+Lemma 3.1.
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of positive real numbers with $c_{k}\geq 1$ for all $k\in\mathbb{N}$.
+If $\mathcal{W}(c_{k})$ is non-empty, then the smallest element in $\mathcal{W}(c_{k})$ exists.
+Proof.
+Since $\mathcal{W}(c_{k})$ is lower bounded and non-empty, $\inf\mathcal{W}(c_{k})$ exists.
+Let $\xi=\inf\mathcal{W}(c_{k})$.
+By the definition of the infimum, there is a sequence $A_{1}\geq A_{2}\geq\cdots$ such that $A_{j}\in\mathcal{W}(c_{k})$ for every $j\in\mathbb{N}$ and $\lim_{j\to\infty}A_{j}=\xi$.
+By the right-side continuity of the floor function, for all fixed $k\in\mathbb{N}$, we have
+$\lfloor\xi^{C_{k}}\rfloor=\lfloor\lim_{j\to\infty}A_{j}^{C_{k}}\rfloor=\lim_{j\to\infty}\lfloor A_{j}^{C_{k}}\rfloor.$
+Thus, we have $\lfloor\xi^{C_{k}}\rfloor\in\mathcal{P}$ since $\lfloor A_{j}^{C_{k}}\rfloor\in\mathcal{P}$ for all $j\in\mathbb{N}$ and $\lim_{j\to\infty}\lfloor A_{j}^{C_{k}}\rfloor$ converges in $\mathbb{Z}$.
+Therefore, $\inf\mathcal{W}(c_{k})=\xi\in\mathcal{W}(c_{k})$, which means that $\xi$ is the smallest element of $\mathcal{W}(c_{k})$.
+∎
+Theorem 3.2.
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers satisfying $c_{1}>0$ and $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$. Then, $\mathcal{W}(c_{k})$ is non-empty.
+Proof.
+See [15, Theorem 3] or [18, Theorem 3.3]. ∎
+Remark 3.3.
+In [15], Matomäki actually showed that $\mathcal{W}(c_{k})$ is uncountable, nowhere dense, and has Lebesgue measure 00 if $c_{k}\geq 2$ for all $k\in\mathbb{N}$.
+Thus, there is a small gap between the initial conditions $c_{1}>0$ in Theorem 3.2 and $c_{1}\geq 2$ in Matomäki’s result, but it is not essential by considering the locally bi-Lipschitz continuous map
+$f_{\alpha}\colon\mathcal{W}(c_{1},c_{2},c_{3},\ldots)\ni A\longmapsto A^{1/\alpha}\in\mathcal{W}(\alpha c_{1},c_{2},c_{3},\ldots)$
+for some fixed real number $\alpha>0$. We refer the readers to [18, Remark 3.4] for more details.
+Corollary 3.4.
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers satisfying $c_{1}>0$ and $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$. Then, the smallest element in $\mathcal{W}(c_{k})$ exists.
+Proof.
+It is clear from Lemma 3.1 and Theorem 3.2. ∎
 Let $b$ be an integer greater than or equal to $3$.
-Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3.
-By [13, Theorem 1.2], the set $\mathcal{W}(c_{k})$ is non-empty, closed, and lower bounded.
-Therefore, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists.
-We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by [13, Theorem 1.3].
-Thus setting $B=\sup_{k\in\mathbb{N}}c_{k}$, we have $B<\infty$.
-Let $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$ which is an infinite set from (3).
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.5.
+By Corollary 3.4, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists.
+We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by Theorem 1.4. Thus, we set
+(3.1) $B=\sup_{k\in\mathbb{N}}c_{k}<\infty,$
+Furthermore, we let
+(3.2) $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$
+which is an infinite set from (3).
 Let $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for all $k\in\mathbb{N}$.
 Let $\theta=21/40$.
-The following Lemmas 3.1 and 3.2 can be seen in [13, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.
+The following Lemmas 3.5 and 3.6 can be seen in [18, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.
 Lemma 3.5.
@@
 By the definition of $\xi$, we see that $p_{k}\leq\xi^{C_{k}}<p_{k}+1$, which implies that $p_{k}^{c_{k+1}}\leq\xi^{C_{k+1}}<(p_{k}+1)^{c_{k+1}}$.
-We recall that $c_{k+1}$ is a positive integer greater than or equal to $2$, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$.
-It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number.
+We recall that $c_{k+1}$ is a positive integer, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$.
+It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number and $c_{k+1}$ is an integer greater than or equal to $2$.
 Therefore, we obtain (3.3).
@@
 Lemma 3.6.
-There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
+Let $\mathcal{I}_{b}$ be as in (3.2). There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
 (3.4) $p_{k}^{c_{k+1}}\leq p_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}.$
-To prove Lemma 3.2, we apply the following result given by Matomäki (see [11, Lemma 9] and [10]).
+Before proving Lemma 3.6, we exhibit the following result given by Matomäki (see [15, Lemma 9] and [14]).
 Theorem 3.7.
-There exist positive constants $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
+There exist positive real numbers $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
 #$\#([n,n+n^{\gamma}]\cap\mathcal{P})\leq d_{1}\frac{n^{\gamma}}{\log n}.$
-By applying this theorem, Matomäki [11] constructed an element in $\mathcal{W}(2)$.
-It is truly astonishing because Mills’ construction tells us that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to concrete an element in $\mathcal{W}(2)$.
+By applying this theorem, Matomäki [15] constructed an element in $\mathcal{W}(2)$ as Theorem 3.2.
+It is truly astonishing because Mills’ construction suggests that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to construct an element in $\mathcal{W}(2)$.
 This problem is strongly connected with Legendre’s conjecture which asserts that for every $n$ there exists a prime number between $n^{2}$ and $(n+1)^{2}$.
@@
 Lemma 3.8.
-Let $c$ be a real number in the interval $[2,B]$, and $\epsilon$ be a sufficiently small positive real number.
-Let $X_{0}$ be a sufficiently large positive parameter depending only on $B$ and $\epsilon$.
-Let $X$ and $\eta$ be real numbers satisfying $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$.
-Suppose that
-(3.3) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X}$
-for some constant $d_{2}>0$. Then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
+Let $E$ and $c$ be real numbers with $2\leq c\leq E$.
+Let $d_{1}$ be as in Theorem 3.7.
+Let $\epsilon\in(0,1)$ and $d_{2}>0$.
+Then, there exists $X_{0}=X_{0}(E,\epsilon,d_{2})>0$ such that for all real numbers $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$, if we have
+(3.5) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X},$
+then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
 (3.6) #$\#([q^{c},q^{c}+q^{c-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c-1}}{\log q^{c}}.$
 Proof.
-Let $R=[X,X+X^{\eta}]\cap\mathcal{P}$. Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$. Indeed, we have
+Let $D$ be as in Theorem 3.7.
+Let $X_{0}$ be a sufficiently large parameter depending only on $E$, $\epsilon$, and $d_{2}$.
+Take arbitrary real numbers $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$.
+We suppose that (3.5) is true.
+Then, let $R=[X,X+X^{\eta}]\cap\mathcal{P}$.
+Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$.
+Indeed, we have
 $X^{c}$ $\leq q^{c}<q^{c}+q^{c-1}=q^{c}(1+q^{-1})\leq(X+X^{\eta})^{c}(1+X^{-1})$
-$\leq X^{c}(1+X^{\eta-1})^{B}(1+X^{-1})\leq 2X^{c}$
-since $\eta\leq 1-\epsilon$ and $X$ is sufficiently large. Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$. By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
-(3.5) $\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
-$\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
+$\leq X^{c}(1+X^{\eta-1})^{E}(1+X^{-1})\leq 2X^{c}$
+since $\eta\leq 1-\epsilon$, $X\geq X_{0}$, and $X_{0}$ is sufficiently large.
+Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$.
+By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
+$\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
+(3.7) $\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
 since $X\geq X_{0}$ and $X_{0}$ is sufficiently large.
@@
 We will find $w\in\mathcal{W}(c_{k})$ such that $w<\xi$ by Matomäki’s construction, which will imply a contradiction to the minimality of $\xi$.
-We now take $k_{0}$ as a sufficiently large positive parameter.
-By the assumption, we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.6).
-By Lemma 3.1, it is false that $p_{k}^{c_{k+1}}>p_{k+1}$, and hence we have
+Let $B$ be as in (3.1), let $\epsilon=\min(1/B,1/4)$, and let $d_{2}=\min(d_{0},d_{1})$, where $d_{0}$ and $d_{1}$ denote as in Theorems 2.1 and 3.7, respectively.
+Let $X_{0}=X_{0}(B,\epsilon,d_{2})$ be as in Lemma 3.8.
+We now take $k_{0}$ as a sufficiently large positive parameter depending only on $B$, $\epsilon$, and $d_{2}$.
+By the assumption (3.8), we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.8).
+By Lemma 3.5, the inequality $p_{k}^{c_{k+1}}>p_{k+1}$ does not hold, and hence we have
 (3.9) $p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}.$
 Then, we set $q_{k}=p_{k}$. Theorem 2.1 with $x\coloneqq q_{k}^{c_{k+1}}$ implies that
-#$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
-By Lemma 3.4 with $X\coloneqq q_{k}^{c_{k+1}}$, $\eta\coloneqq\theta$, and $c\coloneqq c_{k+2}$, there exists $q_{k+1}\in\mathcal{P}$ satisfying
-(3.8) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1})$
-(3.9) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}}.$
+(3.10) #$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
+We may assume that $q_{k}^{c_{k+1}}\geq X_{0}=X_{0}(B,\epsilon,d_{2})$ since $k\geq k_{0}$ and $k_{0}=k_{0}(B,\epsilon,d_{2})$ is sufficiently large. By (3.10) and Lemma 3.8 with
+$E\coloneqq B,\quad c\coloneqq c_{k+2}\in[2,B],\quad\epsilon\coloneqq\epsilon,\quad d_{2}\coloneqq d_{2},$
+$X\coloneqq q_{k}^{c_{k+1}},\quad\eta\coloneqq\theta\in[1/2,3/4]\subseteq[1/2,1-\epsilon],$
+there exists $q_{k+1}\in\mathcal{P}$ satisfying
+(3.11) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1}),$
+(3.12) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}},$
+where the last inequality of (3.11) follows from $c_{k+1}=b\geq 3$ and $\theta=21/40$.
 Let us construct a sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying the following properties: for every $m\geq k$, we have
 (3.13) $q_{m}^{c_{m+1}}\leq q_{m+1}\leq q_{m}^{c_{m+1}}+q_{m}^{c_{m+1}-1},$
-(3.11) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{3}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}},$
-where $d_{3}$ is a small positive constant.
+(3.14) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}}.$
 In the case $m=k$, both (3.13) and (3.14) are true by (3.11) and (3.12).
 We assume that there exists a sequence $(q_{m})_{m=k}^{M}$ of prime numbers satisfying (3.13) and (3.14) for every $m=k,\ldots,M-1$, where $M\geq k+1$.
-Then by applying Lemma 3.4 with $X\coloneqq q_{M}^{c_{M+1}}$, $\eta\coloneqq(c_{M+1}-1)/c_{M+1}$, and $c\coloneqq c_{M+2}$, there exists a prime number $q_{M+1}$ satisfying (3.10) and (3.11) with $m=M$.
+Then by applying (3.14) with $m=M-1$ and Lemma 3.8 with
+$E\coloneqq B,\quad c\coloneqq c_{M+2}\in[2,B],\quad\epsilon\coloneqq\epsilon,\quad d_{2}\coloneqq d_{2},$
+$X\coloneqq q_{M}^{c_{M+1}},\quad\eta\coloneqq(c_{M+1}-1)/c_{M+1}\in[1/2,1-1/B]\subseteq[1/2,1-\epsilon],$
+we find that there exists a prime number $q_{M+1}$ satisfying (3.13) and (3.14) with $m=M$.
 By induction, we obtain an infinite sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying (3.13) and (3.14).
@@
 (3.15) $q_{m}^{1/C_{m}}\leq q_{m+1}^{1/C_{m+1}}<(q_{m+1}+1)^{1/C_{m+1}}<(q_{m}+1)^{1/C_{m}}$
-for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=w$ and $\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}=w^{\prime}$ exist.
+for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}\eqqcolon w$ exist.
 In addition, $w\in\mathcal{W}(c_{k})$ holds by (3.15).
-On the other hand, combining (3.7) and (3.8),
+On the other hand, combining (3.9) and (3.11), we have
 $\lfloor w^{C_{k+1}}\rfloor=q_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}=\lfloor\xi^{C_{k+1}}\rfloor.$
@@
 Therefore, for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$, we have
-(3.16) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{n}(1+p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}=p_{k}+O(p_{k}^{c_{k+1}(\theta-1)+1}).$
+(3.19) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{k}(1+2p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}\leq p_{k}+2p_{k}^{c_{k+1}(\theta-1)+1}.$
 We note that $c_{k+1}(\theta-1)+1=c_{k+1}(\theta-1+\frac{1}{b})=-c_{k+1}\theta_{b}<0$ since $c_{k+1}=b$ and $b\geq 3$. Therefore, we get (3.16). Furthermore, Lemma 3.5 leads to
@@
 Combining (3.16) and (3.20), we obtain (3.17). ∎
-Remark 3.6.
-We can observe the main difficulties of Question 1.4 in (3.16). If $c_{1}=c_{2}=\cdots=2$, then roughly speaking, (3.16) should be
-$p_{k}\leq\xi^{2^{k}}<p_{k}+p_{k}^{2(1/2-1)+1}=p_{k}+1.$
-It does not suffice to obtain (3.14). Further, by Matomäki’s construction, we do not know which specific prime number $q$ satisfies (3.4). This is also one of the main difficulties.
 4. Proof of Theorem 1.5
@@
 Let $b$ be an integer greater than or equal to $3$.
-Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3.
+Let $(c_{k})_{k=1}^{\infty}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.5.
 Let $\xi$ be the smallest element in $\mathcal{W}(c_{k})$.
@@
 We now show the following claim, where $\theta_{b}$ is defined in Lemma 3.9.
-Claim. We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
+Claim: We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
 If $\ell=1$, then $\beta(=\xi^{C_{m}})$ is a positive integer. Therefore,
@@
 Thus, we obtain the claim.
-Case ($b\geq 5$): It follows that
+Case ($b\geq 5$): We now recall that $\theta_{b}=1-\theta-1/b=19/40-1/b$, and so
 $b\theta_{b}\geq\frac{19\cdot 5}{40}-1\geq\frac{19}{8}-1=\frac{11}{8}\quad\therefore\quad(b\theta_{b})^{-1}+1\leq\frac{8}{11}+1<2,$
-which is a contradiction to the claim. Therefore, $\xi$ is transcendental.
+a contradiction to the claim. Therefore, $\xi$ is transcendental.
 Case ($b=4$): By simple calculation,
@@
 Lemma 4.2.
-Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1.
+Let $b$, $(c_{k})_{k=1}^{\infty}$, and $\xi$ be as in Lemma 4.1.
 Suppose that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number.
@@
 Further, we recall that $|\beta_{j}|<1$ for every $2\leq j\leq\ell$, which implies that $\delta_{k}\in(-1/2,1/2)$ since $k$ is sufficiently large.
-If $-1/2<\delta_{k}\leq 0$, then $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$.
 If $20<\delta_{k}<1/$ 2, then by (3.16), $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}$ holds for some constant $K_{0}>0$.
@@
 $\delta_{k}\leq t_{k}-p_{k}\leq\delta_{k}+K_{0}p_{k}^{-c_{k+1}\theta_{b}}.$
-Since $t_{k}-p_{k}\in\mathbb{Z}$ and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough. ∎
+Since we have $\delta_{k}>0$, $t_{k}-p_{k}\in\mathbb{Z}$, and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough.
+Thus, $-1/2<\delta_{k}\leq 0$ holds and $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$.
+∎
 Lemma 4.3.
-Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
+Let $b$, $(c_{k})_{k=1}^{\infty}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
 Proof.
@@
 By applying Lemma 4.2, for sufficiently large $k\in\mathcal{I}_{3}$, we have $p_{k}^{3}=p_{k+1}+3x_{1}x_{2}p_{k}$.
-By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$.
-This is a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
+By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$, a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
 Case ($b=4$): Let $k$ be a sufficiently large integer in $\mathcal{I}_{4}\cap[m,\infty)$.
@@
 $\beta_{1}^{d_{k}}+\beta_{2}^{d_{k}}=p_{k}=\lfloor\beta_{1}^{d_{k}}\rfloor\leq\beta_{1}^{d_{k}}.$
-Then $\beta_{2}^{d_{k}}\leq 0$. This is a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
+Then $\beta_{2}^{d_{k}}\leq 0$, a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
 Combining Lemmas 4.1 and 4.3 completes the proof of Theorem 1.5.
@@
 5. Further discussions
-We will give several assumptions leading to the transcendency of $\xi_{3}$.
+We give several assumptions leading to the transcendency of $\xi_{3}$.
 Proposition 5.1.
@@
 Thus, by Remark 4.4, we have $p_{k+1}\equiv p_{k}\mod 3$ for every sufficiently large $k\geq m$.
-This is a contradiction.
 ∎
@@
 $3=\deg(\xi_{3}^{3^{m}})\leq(3\theta_{3}^{\prime})^{-1}+1=\frac{1}{2-3\theta^{\prime}}+1<\frac{1}{2-3/2}+1=3,$
-which is a contradiction. ∎
+a contradiction. ∎
 Let $p_{n}$ be the $n$th prime number for every $n\in\mathbb{N}$.
 Cramér [7] conjectured that $p_{n+1}-p_{n}=O((\log p_{n})^{2})$.
-If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive constant.
+If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive real number.
 Therefore, assuming Cramér’s conjecture, $\xi_{3}$ is transcendental.
+Question 5.3.
+Assuming the Riemann hypothesis, is Mills’ constant transcendental?
+Carneiro, Milinovich, and, Soundararajan [5, Theorem 1.5] showed that for all real numbers $x\geq 4$, there exists a prime number $p\in\left(x,x+\frac{22}{25}\sqrt{x}\log x\right)$ if the Riemann hypothesis is true.
+However, this does not reach the assumption in Proposition 5.2.
+We need more arithmetic (or algebraic) properties of Mills primes to answer the question affirmatively.
 Acknowledgement
 The author would like to thank Professor Shigeki Akiyama for daily attractive discussions on mathematics.
-The author is grateful to Professor Hajime Kaneko for giving useful ideas on the cases $b=3$ and $4$.
-The author is also grateful to Professor Wataru Takeda for finding errors and mistakes.
-The author was supported by JSPS KAKENHI Grant Numbers JP22J00025 and JP22KJ0375.
+The author is grateful to Professor Hajime Kaneko for giving useful ideas to simplify the proof on the cases $b=3$ and $4$.
+The author is also grateful to Professor Wataru Takeda and the referees for finding errors and mistakes.
+The author was supported by JSPS KAKENHI Grant Numbers JP22J00025, JP22KJ0375, and JP25K17223.